OpenStudy (anonymous):

How can I find the inverse function of f(x) = e^x^3 ??

OpenStudy (anonymous):

I know that I have to switch places of 'x' and 'y' and then solve for 'y'. But I need help to solve for 'y' in this case.

OpenStudy (anonymous):

x=e^y^3

OpenStudy (campbell_st):

so that's a good start \[x = e^{y^3}\] any thoughts on the next step..?

OpenStudy (campbell_st):

hint... it involves logs...

OpenStudy (anonymous):

Right, so do we have to take log on both sides?

OpenStudy (campbell_st):

correct... use natural logs and uses the log law \[\ln(e^a) = a\] so \[\ln(x) = y^3\] can you finish from here..?

OpenStudy (anonymous):

Ohhh got it!! Yeah I can take it from here :D Thank you soooo much! I always get confused when it comes to logs >.<

OpenStudy (campbell_st):

glad to help...

OpenStudy (anonymous):

Thanks :)